OpenStudy (anonymous):

In order to purchase a home, a family borrows $267,000 at 10.8% for 15 years. What is the monthly house payment necessary to amortize the loan? (Do not round until the final answer. Then, round to the nearest cent.) a)$3,394.52 b)$2,403.00 c)$3,001.27 d)$19,108.39

OpenStudy (dean.shyy):

Take a look at this: http://is.gd/mlAt6B

OpenStudy (anonymous):

it looks like c

OpenStudy (amistre64):

does monthly payments assume that the interest is compounded monthly?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

so, assuming monthly periods\[M_1=B(1+\frac i{12})-P\] \[M_2=M_1(1+\frac i{12})-P\\~~~~~~=B(1+\frac i{12})^2-P(1+\frac{i}{12})-P\] to clean things up, let k = 1+i/12 \[M_3=M_2(k)-P\\~~~~~~=B(k)^3-P(k)^2-P(k)-P\\~~~~~~=B(k)^3-P(k^2+k+1)\] \[M_4=M_3(k)-P\\~~~~~~=B(k)^3-P(k^3+k^2+k+1)\] noticing a pattern of the nth month and the exponents \[M_n=M_{n-1}(k)-P\\~~~~~~=B(k)^n-P(k^{n-1}+...+k^3+k^2+k+1)\] those ks attached to the Ps develop into a geometric sum \[M_n=M_{n-1}(k)-P\\~~~~~~=B(k)^n-P(\frac{1-k^n}{1-k})\] letting n = 18*12 months gives us 216 months \[M_{216}=B(k)^{216}-P(\frac{1-k^{216}}{1-k})\] letting M216 = 0 \[0=B(k)^{216}-P(\frac{1-k^{216}}{1-k})\] \[B(k)^{216}=P(\frac{1-k^{216}}{1-k})\] \[\frac{B(k)^{216}(1-k)}{1-k^{216}}=P\]

OpenStudy (anonymous):

does that match up with the c answer? thats what i got with the calculator?

OpenStudy (amistre64):

dunno yet

OpenStudy (amistre64):

phone call, daughter issues :)

OpenStudy (amistre64):

i get 2809, which is close enough to c to warrant variances in methods

OpenStudy (amistre64):

what method did you use?

OpenStudy (anonymous):

i used the calculator that someone provided

OpenStudy (amistre64):

well, then as long as you hit the right buttons, ill assume its correct ....

OpenStudy (amistre64):

http://www.bankrate.com/calculators/mortgages/amortization-calculator.aspx?ec_id=m1024777&ef_id=dCVP-HaRJnMAAACt%3a20130402193615%3as yeah, i keep getting 2808.49 as a payment even when i use on online calculator

OpenStudy (anonymous):

ok, thanks!

OpenStudy (amistre64):

ill get this right sooner or later lol, 12 months times 15 years = 180 periods .... so yeah, 3001 is fine